40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 4 21:41:09 2013, ericq, Update, Green Locking, Green PDH Characterization 
    Reply  Fri Dec 6 12:45:51 2013, ericq, Update, Green Locking, Green PDH Characterization LoopTF.pdfServoTF.pdfMixerOutput.pdfServoOutput.pdf
Message ID: 9447     Entry time: Fri Dec 6 12:45:51 2013     In reply to: 9442
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: Green PDH Characterization 

Yesterday, made a slew of measurements on the X-arm when locked on green. By tweaking the temperature loop offset and the green input PZT pointing, I was able to get the transmitted green to around 1.0. The PDH board gain was set to 4.0. I had trouble making swept sine measurements of the OLTF; changing the excitation amplitude for different frequency ranges would result in discontinuities in the measured TF, and there was only a pretty narrow band around the UGF that seemed to have reasonable coherence.

So, I used the SR785 as a broadband noise generator and measured the TF via dividing the spectra in regions of coherence. Specifically, I used the "pink noise" option of the SR785. I also used a SR560 as a low pass to get enough noise injected into the lower frequency range to be coherent, while not injecting so much into the higher frequencies that the mode hopped while measuring. 

The servo board TF was easily fitted to a 4th order zpk model via VFIT, but I'm having trouble fitting the OLTF. (There is a feature in the servo TF that I didn't fit. This is a feature that Zach saw [ELOG 9537], and attributed to op amp instability) Plots follow. Also, while these need to be calibrated to show the real noise spectrum of the cavity motion, I'm attaching the voltage noise spectra of the error and control signals as a check that electronics/PD noise isn't dominating either signal. 

LoopTF.pdfServoTF.pdf

MixerOutput.pdfServoOutput.pdf

ELOG V3.1.3-